OpenStudy (anonymous):

ok guys I need your help on this one>>> the dimensions of a rectangle are such that its length is 7inches more than it's width. If the length were doubled, and if the width were decreased by 3 inches. The area would be increased by 168 inches sq. What are the length and width of the rectangle? thanks guys

OpenStudy (anonymous):

put x = width, then x +7 is the length and area is \[x(x+7)\]

OpenStudy (anonymous):

ok then it all goes downhill from there lol

OpenStudy (anonymous):

double the length, incraese the width by 3, get \[(x+3)\times 2(x+7)\]

OpenStudy (anonymous):

and so you know that \[x(x+7)+168=2(x+3)(x+7)\]

OpenStudy (anonymous):

now we solve the quadratic equation

OpenStudy (anonymous):

ok x^2 +7x +168=2x+6+x+7

OpenStudy (anonymous):

except i get a really ugly solution, so maybe i made a mistake

OpenStudy (anonymous):

lol well join the club

OpenStudy (anonymous):

oh because it said "decreased" if only i could read

OpenStudy (anonymous):

\[x(x+7)+168=2(x-3)(x+7)\]

OpenStudy (anonymous):

im getting all kinds of wierd answers on this one I cant match it on my check equation

OpenStudy (anonymous):

\[x (x+7)+168 = 2 x^2+8 x-42\] \[x^2+7 x+168 = 2 x^2+8 x-42\] \[x^2+x-210 = 0\] \[(x+15)(x-14)=0\] \[x=14\]

OpenStudy (anonymous):

are you for certain certain?????????????????????????

OpenStudy (anonymous):

yes it matches mine. thanks